A scenic road around a lake Rotorua is 11 miles in length. Tom leaves the hotel on his bike averaging 10 mph and heads west around the lake. Bill leaves at the same time and heads in the opposite direction at 45 mph. How far from the hotel will Tom have traveled when they meet.
Answer by stanbon(75887)

A scenic road around a lake Rotorua is 11 miles in length.
Tom leaves the hotel on his bike averaging 10 mph and heads west around the lake.
Bill leaves at the same time and heads in the opposite direction at 45 mph.
How far from the hotel will Tom have traveled when they meet.
--------------
Tom DATA
rate = 10 mph ; time = x hrs ; distance = 10x miles
---
Bill DATA:
rate = 45 mph ; time = x hrs ; distance = 45x miles
-----
distance + distance = 11 miles
10x + 45x = 11 miles
55x = 11 miles
x = 1/5 hr
---
Tom's distance = 10*(1/5) = 2 miles
=========================================
cheers,
Stan H.
=============